OpenStudy (anonymous):

Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = x2 - 14

OpenStudy (anonymous):

Shift the graph of y = x2 down 14 units. this is what i got

OpenStudy (phi):

yes btw it would be better to write y= x^2-14

OpenStudy (anonymous):

ur complete right thank you soooo much !!

OpenStudy (phi):

you can check your answer: y=x^2 at x=0 gives y=0*0= 0 y= x^2-14 at x=0 gives 0*0-14= -14 which is down 14 from the first equation. so that is the right answer.